OpenStudy (perl):

Oil consumption is increasing at a rate of 1.9% per year. What is its doubling time? I'm not sure if this is linear increase or exponential. if its linear

OpenStudy (perl):

, no , this website says its exponential http://www.algebra.com/algebra/homework/logarithm/logarithm.faq.question.112025.html

OpenStudy (perl):

, if its linear then y = Ao ( 1 + .06*t ) if its exponential then y = Ao ( 1 + .06)^t

OpenStudy (anonymous):

i think its exponential 2 = 1 ( 1 + 0.06)^t now calculate t

OpenStudy (anonymous):

Oh, wait. That's right. I apologize. I graphed it wrong.

OpenStudy (perl):

how will i know if its exponential or not?

OpenStudy (perl):

here is another example http://www.physics.uoguelph.ca/tutorials/exp/intro.html

OpenStudy (perl):

why isnt the following example linear ? "For example, let's look at a population of wee beasties which increases by 10% per year. If there were 100 wee beasties now, there would be an increase of 10 wee beasties after a year. "

OpenStudy (anonymous):

see in question it increases 1.9% per year first year let consumption = x next year = 1.9*x/100 later year = 1.9(1.9*x/100)/100 if you plot its graph you get exponential curve try it

OpenStudy (anonymous):

in your second example population first year = 100 population second year = 110 population third year = 121 it also not linear

OpenStudy (perl):

you need a plus somewhere

OpenStudy (perl):

it should be next year consumption = last year consumption* ( 1 + 1.9/100)

OpenStudy (anonymous):

yes you are write

OpenStudy (perl):

so we have x1 = x x2 = x1 (1.019) x3 = x2 (1.019) = [x1 (1.019) (1.019) ]

OpenStudy (perl):

better start with zero, x0 = x x1 = x0 (1.019) x2 = x1 (1.019) = [x0 (1.019) ] (1.019) = x0 (1.019) ^2

OpenStudy (anonymous):

calculate x1,x2 ,x3 and try to find x2 - x1 = x3 -x2 if its write then only its linear

OpenStudy (perl):

so did you agree with my math?

OpenStudy (perl):

right, clearly x2-x1 != x3 - x2

OpenStudy (anonymous):

yes i agree so what you think its linear or exponential

OpenStudy (perl):

well... exponential

OpenStudy (perl):

i think its a little ambiguous however

OpenStudy (perl):

, you could also have y = x0 ( 1 + .019 * t ) , no ?

OpenStudy (perl):

ok how would you word it so that it is linear?

OpenStudy (anonymous):

let if increase every year is same then its linear for example if population of city inc by 20 people every year then we can say that its linear

OpenStudy (perl):

ok , so we wouldnt use percentage for a linear increase

OpenStudy (anonymous):

yes

OpenStudy (perl):

if there is a percentage growth , like population percentage growth , grows at a constant percentage, it is exponential growth

OpenStudy (anonymous):

yes

OpenStudy (perl):

got it,

OpenStudy (anonymous):

ok then

OpenStudy (perl):

ok now doubling time makes sense

OpenStudy (perl):

doubling time is the time it takes for the initial amount to grow 100% , leaving you with double

OpenStudy (anonymous):

yes did you get your answer?

OpenStudy (perl):

yes, the math i didnt have trouble with. just the understanding

OpenStudy (perl):

i have a much harder problem, differential equation

OpenStudy (perl):

are you ready?

OpenStudy (perl):

its a logistic growth problem

OpenStudy (anonymous):

ok then now you can close the qustion after giving me medal...............lol

OpenStudy (perl):

you stuck it out , good work

OpenStudy (perl):

the rest bailed